# Yuge Tu1, Pengkai Li2, Lei Shi1, Zhiyuan Liu1, Maosong Sun1

### v1

@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Achiam et al. (2023) and Large Multimodal Models (LMMs) Team et al. (2023) present inspiring capabilities in understanding and generating human-like text and realistic images. However, their direct application in embodied AI, where agents interact in physical or simulated environments, is still primitive. LLMs and LMMs lack the necessary grounding Harnad (1990) in physical interactions to operate in these settings effectively.
@@
To achieve generalizable embodied intelligence, two key factors are crucial: language grounding to utilize the extensive knowledge in LMMs, and the expansion of training data for embodied AI. There have been noteworthy efforts in combining embodied AI with LMMs Reed et al. (2022); Brohan et al. (2023). They collect large-scale training data from embodied scenes and train end-to-end models that interpret both language and visual inputs and perform corresponding actions. However, the lack of open-source access to these environments and datasets restricts open-source community-wide progress in this field. Therefore, the academic community urgently requires an open-source platform that facilitates the integration of language grounding in embodied environments and schemes to generate large-scale training data for embodied agents based on LLMs and LMMs.
Towards this aspiration, we introduce LEGENT, an open-source and user-friendly platform that enables scalable training of embodied agents based on LLMs and LMMs. LEGENT contains two parts. First, it provides a 3D embodied environment with the following features: (1) Diverse, realistic, and interactive scenes; (2) Human-like agents with egocentric vision capable of executing actions and engaging in direct language interaction with users; (3) User-friendly interface offering comprehensive support for researchers unfamiliar with 3D environments. Second, LEGENT builds a systematic data generation pipeline for both scene generation and agent behavior, incorporating state-of-the-art algorithms for scene creation Deitke et al. (2022); Yang et al. (2023b) and optimal trajectory generation. In this way, extensive and diverse trajectories of agent behavior with egocentric visual observations and corresponding actions can be generated at scale for embodied agent training.
@@ 2 Related Work
Embodied Environment. Embodied environments are extensively utilized in games Oh et al. (2016) and robotics Kolve et al. (2017); Yan et al. (2018); Xia et al. (2018); Gan et al. (2020); Li et al. (2021); Puig et al. (2023a), with a primary focus on visual AI and reinforcement learning. Some platform focuses on specific embodied tasks, such as manipulation Yu et al. (2020); Makoviychuk et al. (2021), navigation Chang et al. (2017); Dosovitskiy et al. (2017), or planning-oriented agents Puig et al. (2018); Shridhar et al. (2020); Wang et al. (2022). However, the environment setups and data frameworks of existing platforms fall short in accommodating the training of LMMs, which excel in the supervised learning paradigm and necessitate diverse and large-scale data to integrate embodied capability.
LMMs-based Embodied Agent. Noteworthy studies have concentrated on developing embodied models capable of end-to-end operation, as demonstrated in the works of Reed et al. (2022); Brohan et al. (2023); Belkhale et al. (2024). Despite being generalizable, the datasets and models in these studies are not publicly available.
@@
Agent Trajectory Generation. Some research focuses on crafting reward functions to guide small policy models Yu et al. (2023); Xian et al. (2023); Wang et al. (2023); Ma et al. (2023). However, there will be huge costs and instability when applying reward-based training to large foundation models. Meanwhile, pioneering efforts have been made in code generation for robotics Liang et al. (2023); Singh et al. (2023); Vemprala et al. (2023); Huang et al. (2023) and trajectory generation for imitation learning Kamath et al. (2023). We follow this paradigm and further scale agent trajectory generation to cover a wide range of tasks.
@@ 3.1 Scene
Scalable Assets. LEGENT supports importing customized objects, including the user-supplied 3D objects, objects from existing datasets Deitke et al. (2023) and those created by generative models Siddiqui et al. (2023); Wang et al. (2024). We choose glTF as the import format for its openness and broad compatibility. This feature grants users the flexibility to customize the scene by strategically placing these assets or integrating them seamlessly into scene generation algorithms.
@@ 3.2 Agent
Realistic Animation. LEGENT features precise humanoid animations using inverse kinematics and spatial algorithms, enabling lifelike movements, which is important to enhance nonverbal interactions in AI systems and contribute to robotic control and text-to-motion research. Also, when combined with egocentric vision, it offers a cost-effective alternative for immersive experiences similar to the Ego4D Grauman et al. (2022), which requires a huge cost to collect at scale.
@@ 4.1 Scene Generation
Figure 2: Generated scenes by ProcTHOR (Left) and Holodeck (Right).
@@ 4.2 Task Generation
Scene Generation for Given Tasks. This approach efficiently generates large-scale samples for specific tasks, provided that the scene generation algorithm fulfills the required criteria. For instance, when the task involves querying an object’s location, the algorithm generates a scene that includes the object and its receptacle, inherently creating question-answering annotations. As shown in Table 2, we provide some basic task templates that are ideal for creating large-scale scenes, which are particularly useful for pretraining fundamental capabilities of embodied control, spatial comprehension, and basic language grounding across diverse scenes.
@@ 4.3 Trajectory Generation
We use LLMs and controllers to label the ground truth actions. Inspired by pioneering works in code generation for robotics, we utilize LLMs to write intermediate codes from provided state descriptions and instructions. These codes are instantiated as multi-step controllers, designed to calculate the optimal actions at each step given the internal states of the environment. Each controller operates in a step-by-step manner in the environment, with visual observations collected during the process.
@@
Note that the code-writing is annotation-oriented. Even though LLMs can directly answer the question from the state description, it still invokes “find”. Then the code “find” is instantiated as a multi-step controller that utilizes pathfinding algorithms Hart et al. (1968) incorporating visibility checks. The pathfinding algorithm calculates the waypoints of the shortest path from the agent to the target object using a navigation mesh. The controller then calculates the controls of the agent to navigate along these waypoints. For instance, in the first observation shown in Figure 3, the agent needs to rotate 59 degrees to the left to orient to the next waypoint, resulting in the action “rotate_right(-59)”. Similarly, in the second observation, the agent needs to perform certain actions to move to the subsequent waypoint. This multi-step controller concludes when the target object enters the agent’s field of view. LEGENT records the visual observations and actions during this process as a trajectory, which can be exported as a video or an image-text interleaved sequence. The actions use a unified code representation, compatible with the outputs of LMMs.
@@ 4.4 Prototype Experiments
We conduct a prototype experiment to assess the utility of generated data on two embodied tasks: “Come Here” for social navigation Puig et al. (2023a) and “Where Is” for embodied question answering Das et al. (2018). Task complexity varied from navigating in one room to the more intricate two rooms. We generate 1k and 10k trajectories for the initial three tasks (“Come Here” in one or two rooms and “Where Is” in one room) and assess the models on 100 trajectories across all four tasks. The “Where Is” task in the two-room setting serves as a generalization test, which is not included in the training data.
Due to the lack of powerful video understanding models, we temporarily only focus on the observation at the end of each continuous action, formulating one trajectory as an image-text interleaved sequence. We utilize VILA-7B Lin et al. (2023) as our backbone due to its capability in interleaved inputs. We train the model to output current action based on task descriptions and interleaved context of previous observations and actions.
@@ 4.5 Demo of LEGENT
The demo video of LEGENT is available at the link [^1] https://video.legent.ai, which is partially shown in Figure LABEL:fig:crown. The demonstration exemplifies the engagement with embodied agents in LEGENT, primarily leveraging LLMs and controllers described in Section 4.3. With advancements in LMMs’ capability of egocentric perception and control, we foresee the evolution of this demonstration into a fully embodied experience, independent of any extra internal information. We will also pursue this goal by further scaling the data generation for model training.
@@ 5 Conclusion and Future Work
Limitations
In the context of imitation learning, the persistent challenge of insufficient exploration and handling out-of-distribution inputs during inference underscores the need for further enhancements and strategies within the data generation pipeline, a component that is currently not integrated into our system. Furthermore, large-scale experiments have yet to be conducted. We leave this to the near future work.
Ethics Statement
The development of LEGENT prioritizes ethical considerations across all aspects of its use and implementation. We uphold data privacy and security, ensuring compliance with relevant data protection laws. We strictly adhere to legal standards and encourage ethical use of our platform. We are committed to continuous evaluation of the ethical implications of our work and engaging with the community to address emerging concerns and ensure a positive impact on society.

### v2

# Yuge Tu, Pengkai Li, Lei Shi, Zhiyuan Liu, Maosong Sun

@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023a, b) and Large Multimodal Models (LMMs) OpenAI (2023); Team et al. (2023); Liu et al. (2024); Hu et al. (2024) present inspiring capabilities in understanding and generating human-like text and realistic images. However, their direct application in embodied AI, where agents interact in physical or simulated environments, is still primitive. LLMs and LMMs lack the necessary grounding Harnad (1990) in physical interactions to operate in these settings effectively.
@@
To achieve generalizable embodied intelligence, two key factors are crucial: language grounding to utilize the extensive knowledge in LMMs, and the expansion of training data for embodied AI. There have been noteworthy efforts in combining embodied AI with LMMs Reed et al. (2022); Brohan et al. (2023). They collect large-scale training data from embodied scenes and train end-to-end models that interpret both language and visual inputs and perform corresponding actions. However, the lack of open-source access to these environments and datasets restricts open-source community-wide progress in this field. Therefore, the academic community urgently requires an open platform that facilitates the integration of language grounding in embodied environments and schemes to generate large-scale training data for embodied agents based on LLMs and LMMs.
Towards this aspiration, we introduce LEGENT, an open and user-friendly platform that enables scalable training of embodied agents based on LLMs and LMMs. LEGENT contains two parts. First, it provides a 3D embodied environment with the following features: (1) Diverse, realistic, and interactive scenes; (2) Human-like agents with egocentric vision capable of executing actions and engaging in direct language interaction with users; (3) User-friendly interface offering comprehensive support for researchers unfamiliar with 3D environments. Second, LEGENT builds a systematic data generation pipeline for both scene generation and agent behavior, incorporating state-of-the-art algorithms for scene creation Deitke et al. (2022); Yang et al. (2023b) and trajectory generation. In this way, extensive and diverse trajectories of agent behavior with egocentric visual observations and corresponding actions can be generated at scale for embodied agent training.
@@ 2 Related Work
Embodied Environment. Embodied environments are extensively utilized in games Johnson et al. (2016); Oh et al. (2016); Beattie et al. (2016) and robotics Kolve et al. (2017); Yan et al. (2018); Xia et al. (2018); Gan et al. (2020); Li et al. (2021); Puig et al. (2023a), with a primary focus on visual AI and reinforcement learning. Some platform focuses on specific embodied tasks, such as manipulation Yu et al. (2020); Makoviychuk et al. (2021), navigation Chang et al. (2017); Dosovitskiy et al. (2017), or planning-oriented agents Puig et al. (2018); Shridhar et al. (2020); Wang et al. (2022). However, the environment setups and data frameworks of existing platforms fall short in accommodating the training of LMMs. LMMs excel in the supervised learning paradigm and necessitate diverse and large-scale data to integrate embodied capability. Existing platforms are not yet ready to scale, including: the primarily supported reinforcement learning methods require careful reward engineering, the diversity of the training data cannot be easily expanded, and collecting data for imitation learning on these platforms requires manual effort.
LMMs-based Embodied Agent. Noteworthy studies have concentrated on developing embodied models capable of end-to-end operation, as demonstrated in the works of  Reed et al. (2022); Brohan et al. (2023); Belkhale et al. (2024). However, the datasets and models in these studies are not publicly available.
@@
Agent Trajectory Generation. Some research focuses on crafting reward functions to guide small policy models Yu et al. (2023); Xian et al. (2023); Wang et al. (2023); Ma et al. (2023). However, there will be huge costs and instability when applying reward-based training to large foundation models. Meanwhile, pioneering efforts have been made in code generation for robotics Liang et al. (2023); Singh et al. (2023); Vemprala et al. (2023); Huang et al. (2023) and trajectory generation for imitation learning Garrett et al. (2021); Kamath et al. (2023); Dalal et al. (2023). These efforts align with our approach to generating large-scale embodied trajectories for training LMMs.
@@ 3.1 Scene
Scalable Assets. LEGENT supports importing customized objects at runtime, including user-supplied 3D objects, objects from existing datasets Deitke et al. (2023) and those created by generative models Siddiqui et al. (2023); Wang et al. (2024), as illustrated in Fig. 2. We choose glTF as the import format for its openness and broad compatibility. This feature grants users the flexibility to customize the scene by strategically placing these assets or integrating them seamlessly into scene generation algorithms.
Figure 2: Examples of importing external assets: user-supplied assets (left); existing datasets (middle); assets generated by generative models (right).
@@ 3.2 Agent
Realistic Animation. LEGENT features precise humanoid animations using inverse kinematics and spatial algorithms, enabling lifelike movements, as shown in Fig. 3. It is important for enhancing nonverbal interactions in AI systems and contributes to robotic control and text-to-motion research. Also, when combined with egocentric vision, it offers a cost-effective alternative for immersive experiences similar to Ego4D Grauman et al. (2022), which requires a huge cost to collect.
Figure 3: An example of humanoid animations, demonstrating accurate object grasping and body movement through spatial planning and inverse kinematics.
@@ 4.1 Scene Generation
Figure 4: Examples of generated scenes.
@@ 4.2 Task Generation
Scene Generation for Given Tasks. This approach efficiently generates large-scale samples for specific tasks based on the scene generation algorithm. For instance, when the task involves querying an object’s location, the algorithm generates a scene that includes the object and its receptacle, inherently creating question-answering annotations. As shown in Table 2, we provide some basic task templates that are ideal for creating large-scale scenes, which are particularly useful for pretraining fundamental capabilities of embodied control, spatial comprehension, and basic language grounding across diverse scenes.
@@ 4.3 Trajectory Generation
We use LLMs and motion planners to label the ground truth actions. Inspired by pioneering works in code generation for robotics, we utilize LLMs to write intermediate codes from provided state descriptions and instructions. These codes are instantiated as motion planners, designed to calculate the optimal actions at each step given the internal states of the environment. Each motion planner operates in a step-by-step manner in the environment, with visual observations collected during the process. This approach is consistent with the concept of Task and Motion Planning (TAMP) Garrett et al. (2021) in robotics, where the LLMs and the motion planners respectively fulfill the functions of task planning and motion planning.
@@
Note that the code-writing is annotation-oriented. Even though LLMs can directly answer the question from the state description, it still invokes “find”. Then the code “find” is instantiated as a motion planner that utilizes pathfinding algorithms Hart et al. (1968) incorporating visibility checks. The pathfinding algorithm calculates the waypoints of the shortest path from the agent to the target object using a navigation mesh. The motion planner then calculates the controls of the agent to navigate along these waypoints. For instance, in the first observation shown in Figure 5, the agent needs to rotate 59 degrees to the left to orient to the next waypoint, resulting in the action “rotate_right(-59)”. Similarly, in the second observation, the agent needs to perform certain actions to move to the subsequent waypoint. This motion planner concludes when the target object enters the agent’s field of view. LEGENT records visual observations and actions during this process as a trajectory, which can be exported as a video or an image-text interleaved sequence. The actions use a unified code representation, compatible with the outputs of LMMs.
@@ 4.4 Prototype Experiments
Figure 6: The vision-language-action model architecture used in the prototype experiments.
We conduct a prototype experiment to assess the utility of generated data on two embodied tasks: “Come Here” for navigation and “Where Is” for embodied question answering Das et al. (2018). Task complexity varied from navigating in one room to the more intricate two rooms. We generate 1k and 10k trajectories for the initial three tasks (“Come Here” in one or two rooms and “Where Is” in one room) and assess the models on 100 trajectories across all four tasks. The “Where Is” task in the two-room setting serves as a generalization test, which is not included in the training data.
Due to the lack of powerful video understanding models, we temporarily only use the observation at the end of each continuous action, formulating one trajectory as an image-text interleaved sequence. We utilize VILA-7B Lin et al. (2024) as our backbone due to its capability in interleaved inputs. As illustrated in Fig. 6, we train the vision-language-action (VLA) model to predict current action based on task descriptions and interleaved context of previous observations and actions,.
@@ 4.5 Demo of LEGENT
The demo video of LEGENT is available at the link [^1] https://video.legent.ai, which is partially shown in Fig. LABEL:fig:crown. The demonstration exemplifies the engagement with embodied agents in LEGENT, primarily leveraging LLMs and motion planners described in Section 4.3. With advancements in LMMs’ capability of egocentric perception and control, we foresee the evolution of this demonstration into a fully embodied experience, independent of any extra internal information. We will also pursue this goal by further scaling the data generation for model training.